OpenStudy (anonymous):

an object starts from rest and falls freely. What is the velocity of the object at the end of 9.00 seconds?

OpenStudy (anonymous):

V@9sec = g * 9 where g = 9.8 m/sec (acceleration due to gravity) from the relation : v=u + a*t

OpenStudy (espex):

\[v_f = v_i + at\] \[v_f = 0 + 9.8\frac{m}s \times 9s\]

OpenStudy (anonymous):

88.2 m/s2

OpenStudy (anonymous):

Well, technically it would be 88.2 m/s... Because it's 9.8 m/s^2 * 9 seconds... so one of the seconds cancels out, leaving the 88.2 m/s.